# Data-Driven Parametrization of Molecular Mechanics Force Fields for Expansive Chemical Space Coverage

### v1

@@ 2.1 Molecular Mechanics Force Field
$\begin{split}E^{\mathrm{MM}}=&E_{\mathrm{bonded}}^{\mathrm{MM}}+E_{\mathrm{non-bonded}}^{\mathrm{MM}}\\ E_{\mathrm{bonded}}^{\mathrm{MM}}=&E_{\mathrm{bond}}^{\mathrm{MM}}+E_{\mathrm{angle}}^{\mathrm{MM}}+E_{\mathrm{proper}}^{\mathrm{MM}}+E_{\mathrm{improper}}^{\mathrm{MM}}\\ =&\sum_{\mathrm{bonds}}\frac{1}{2}k_{r,ij}(r_{ij}-r_{ij}^{0})^{2}+\sum_{\mathrm{angles}}\frac{1}{2}k_{\theta,ijk}(\theta_{ijk}-\theta_{ijk}^{0})^{2}\\ &+\sum_{\mathrm{propers}}\sum_{n_{\phi}}k_{\phi,ijkl}^{n_{\phi}}\left[1+\cos(n_{\phi}\phi_{ijkl}-\phi_{ijkl}^{n_{\phi},0})\right]\\ &+\sum_{\mathrm{impropers}}\sum_{n_{\psi}}k_{\psi,ijkl}^{n_{\psi}}\left[1+\cos(n_{\psi}\psi_{ijkl}-\psi_{ijkl}^{n_{\psi},0})\right]\\ E_{\mathrm{non-bonded}}^{\mathrm{MM}}=&E_{\mathrm{vdW}}^{\mathrm{MM}}+E_{\mathrm{Coulomb}}^{\mathrm{MM}}\\ =&\sum_{i<j}\left[4\epsilon_{ij}\left(\frac{\sigma_{ij}^{12}}{r_{ij}^{12}}-\frac{\sigma_{ij}^{6}}{r_{ij}^{6}}\right)+\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r_{ij}}\right]\end{split}$ (1)
where the bond lengths $r$, angles $\theta$, proper torsion angles $\phi$ and improper torsion angles $\psi$ are rotationally invariant internal coordinates. The force field parameters include bonded parameters (equilibrium values $r^{0}$, $\theta^{0}$, $\phi^{0}$ and $\psi^{0}$, and force constants $k_{r}$, $k_{\theta}$, $k_{\phi}$ and $k_{\psi}$), and non-bonded parameters (van der Waals (vdW) parameters $\sigma$ and $\epsilon$, and partial charges $q$). A well-trained transferrable MMFF shall predict these parameters accurately for any given molecules. In this article, “torsion” refers to proper torsion unless otherwise specified. We also fix the phase angles $\phi_{ijkl}^{n_{\phi},0}$ and $\psi_{ijkl}^{n_{\psi},0}$ at 0 for odd $n$ and 180 for even $n$, and therefore ensure that the torsional energy is independent of the order being $ijkl$ or $lkji$.
@@
Additionally, there are two key philosophies worth considering when building general small molecule MMFFs: (1) Force field parameters should be dominated by local structures, so that the parameters trained from small molecules can be consistently transferred to similar structures in relatively large molecules. 9, 11 (2) Torsional energy profiles should be accurately captured, because the quality of torsion parameters strongly affects the protein-ligand binding affinity predictions. 32, 11, 33
@@ 2.2.2 Quantum chemistry methods and workflow
From the 2.4 million fragments, we created two QM datasets, namely the optimization dataset and the torsion dataset, at the B3LYP-D3(BJ)/DZVP level of theory. The same QM method is also used by OpenFF to generate their training data. 18, 19 This method achieves a good balance between accuracy (relative to CCSD(T)/CBS) and computational cost in recent benchmarks. 38, 39 For molecular conformational PES, more advanced methods like $\omega$B97M-v are significantly more expensive but only marginally more accurate.
The optimization dataset was generated from the entire 2.4 million fragments. The 3D conformations of each fragment was initially generated by RDKit from its SMILES string, and then optimized using the geomeTRIC 40 optimizer at the chosen QM level. The relaxed structure is verified where no accidental bond breaking or formation happened during the structural relaxation. Finally, the Hessian matrix is calculated for each fragment using Q-Chem 641. The hessian matrix is further verified by checking the six smallest eigenvalues to ensure that a true local minimum is captured. More details about the workflow and the filtering criteria are provided in the supporting information.
The torsion dataset consists of two subsets, non-ring torsions and in-ring torsions, curated separately. For each non-ring torsion fragment, we first rotate the torsion angle in 15 increments from the optimized conformation, creating 24 initial frames that were then separately relaxed using the geomeTRIC optimizer with the rotated torsion angle constrained. While the in-ring torsions were scanned using a sequential frame-by-frame approach, with an early stopping strategy if the conformational energy is more than 20 kcal/mol higher than the unconstrained relaxed conformation, since those high-energy regimes are irrelevant in room-temperature molecular dynamics simulations. For both non-ring and in-ring subsets, the resultant conformers were also carefully filtered by the bond breaking or formation criteria. Due to the high computational cost of torsion scan, these calculations were performed with GPU4PySCF42 for significant acceleration and cost reduction. The results were also verified to be consistent with Q-Chem.
@@ 2.3 ByteFF Model
ByteFF predicts both the bonded and non-bonded force field terms in one-pass. It leverages the state-of-the-art Edge-augmented Graph Transformer (EGT) 43 as its backbone to achieve accuracy and transferability. The output of EGT is further processed by a series of MLPs to derive force field parameters while preserving both the total charge and the physical symmetry of each force field term. The ML model structure is shown in Fig. 1 and consists of three major parts: featurization, GNN, and output layers. They are briefly described here, and more details are given in the supporting information.
Figure 1: Model structure of ByteFF. ByteFF predicts MMFF parameters in three steps. First, atom and bond features are extracted from the molecular graph and then projected into embeddings. Then, an edge-augmented graph transformer (EGT)43 is used to synergize the edge embeddings with the node-based attention mechanism. Lastly, the output MLPs derive force field parameters while preserving the molecular symmetry and total charge.
Featurization layers. For each molecule, we first convert the molecule structure into a graph, and detect chemically equivalent atoms and bonds with bond orders ignored. Atomic features are then detected from this graph, including the element type, formal charge, ring connectivity and minimum ring size. Additionally, to better describe the chemical environment of each atom, bond features including bond order and whether the bond is in ring, are also extracted. Replacing the discrete atom-typing rules used in conventional MMFFs, atom and bond features are projected into continuous vectorial embedding spaces, and concatenated into node and edge feature embeddings, respectively. Importantly, embeddings of chemically equivalent atoms and bonds are averaged before passed to the next layers, which helps to preserve the chemical symmetry of the molecule.
GNN layers. In the GNN layers, a modified EGT architecture is employed to learn the chemical environment of molecules from both atom and bond features. This architecture synergizes the node and edge features in the attention mechanism to efficiently capture the structural information of the molecule. To speed up training and inference, the original global self-attention in EGT is localized such that attention is applied only to a local neighborhood of each node.
Output module. After a few layers of localized EGT, atom and bond hidden states generated by GNN layers are passed to the output layers to predict each force field term individually. For each bonded term, the corresponding output block is designed to preserve the necessary symmetry, respectively. For the partial charge term, the output block is designed to preserve the total charge of the molecule following a bond-charge-correction (BCC) style postprocessing.
Additionally, an ensemble of three models randomly initialized is trained for improved predictive performance and uncertainty quantification.
@@ 2.4 Training Procedures
In the pre-training stage, the non-bonded parameters and force constants of proper torsions were fitted to GAFF2 with mean squared error (MSE) losses, while the equilibrium values of bonded parameters were fitted with energy-based loss functions, using the optimization dataset. Meanwhile, the force constants of bonded terms in Eq. 1 were trained using the partial Hessian loss, evaluated by the mean absolute percentage error (MAPE). The design of the partial Hessian loss combines the training accuracy with the computational efficiency, which takes the form of Hessian matrices but only fit the blocks corresponding to the target bonds, angles and improper torsions, namely partial Hessian blocks.
Given the significance of torsion profiles in the quality of force fields, in the training stage, we incorporated the curated torsion dataset to fit the force constants of proper torsions, while the other parameters were trained using optimization dataset following the same manner in the pre-training stage. Here, we employed the Boltzmann MSE loss functions 44 into an iterative optimization-and-training process to train the force constants of proper torsions using the torsion dataset. In each iteration, the QM-optimized conformations were refined by the force field, with the torsion angle constrained and atoms restrained. The parameters were then trained on both the optimization and torsion datasets. As the force field’s accuracy improved, the positional restraint force constant was gradually reduced. Additionally, L1-norm regularization was applied to the force constants of proper torsions to restrain redundant degrees of freedom. The combination of pre-training and training stages yields the ByteFF-gopt model.
@@
Datasets and fitting targets optimization dataset: - GAFF $\sigma/\epsilon/q/k_{\phi}$ $E^{\mathrm{MM}}_{\mathrm{bond}}$ / $E^{\mathrm{MM}}_{\mathrm{angle}}$ / $E^{\mathrm{MM}}_{\mathrm{improper}}$ QM Hessian matrix optimization dataset: - GAFF $\sigma/\epsilon/q$ $E^{\mathrm{MM}}_{\mathrm{bond}}$ / $E^{\mathrm{MM}}_{\mathrm{angle}}$ / $E^{\mathrm{MM}}_{\mathrm{improper}}$ Hessian matrix torsion dataset: - QM energy - L1-norm of $k_{\phi}$ optimization dataset: same as training torsion dataset: same as training off-equilibrium dataset: - QM energy/force
@@
Table 1: Summary of training settings for different stages
@@ 2.5 Benchmark Datasets and Metrics
To be compatible with the Amber family of protein force fields, the non-bonded terms of ByteFF are trained to reproduce the am1bcc charges and GAFF vdW parameters. Therefore, in this work we focus on the intramolecular PES performance relative to QM references of ByteFF, and leave the further improvement of non-bonded terms as a future work.
Three different types of benchmarks are used to provide a thorough assessment of ByteFF. First, the large-scale industrial collaborative OpenFFBenchmark public dataset 45 is used to quantify how well the force field-relaxed geometries and energies reproduce the QM-relaxed references, without constraints. Three different metrics are used, namely root-mean-square deviation of atomic positions (RMSD), torsion fingerprint deviation (TFD) 46, and relative energy difference ($\Delta\Delta E$) defined by OpenFF team.47 Second, the crucial torsional energy profile performance is benchmarked on two different datasets, namely the publicly available TorsionNet500 dataset 48 and the BDTorsion dataset curated in-house. For consistency, energies of conformations in TorsionNet500 48 dataset were re-calculated with the B3LYP-D3(BJ)/DZVP level of theory. On these datasets, the root mean squared error (RMSE) and Boltzmann-weighted RMSE of each torsional energy profile are calculated and compared.
Beyond the constrained (torsional energy) and unconstrained (OpenFFBenchmark) optimization benchmarks, the performance of ByteFF is further benchmarked on the QM references energies and forces calculated on the off-equilibrium dataset, which includes subsets of the SPICE dataset49 and the RNA Structure Atlas50. On this dataset, RMSE of energies and forces were calculated to quantify if the PES is properly captured beyond local minima.
More details of these benchmarks can be found in supporting information.
@@ 3.1 Diversity of Various Datasets
Using the Morgan-based torsional fingerprint defined above, we quantify the diversity of SPICE, GEOM and BDTorsion datasets in terms of torsional local environment. Specifically, in all three datasets, only torsions with circular standard deviation 51 greater than 0.3 are included and used in fingerprint analysis. The diversity of each dataset was then visualized using the standard t-SNE (t-Distributed Stochastic Neighbor Embedding) algorithm. In Fig. 2, the scatter dots are colored by the element types of the two center atoms (the rotatable bond) in each torsion. The most frequent rotatable bond is C-C, followed by C-N rotatable bonds and other bonds including C-O, C-S, etc. It is evident that our torsion dataset excels in the diversity and comprehensiveness of its chemical space coverage, while the SPICE and GEOM datasets are more sparse. In comparison, both SPICE and GEOM datasets show obvious vacancies, corresponding to the absence of certain chemical torsion patterns. For example, the uncovered areas around (-45, -20) in Fig. 2 (a) and (b) correspond to in-ring torsions of protonated pyridine and imidazole, as well as their derivatives, which are indispensable for the training of force fields for bio-organic molecules. Such an expansive coverage of the chemical space in the training dataset provides a solid foundation for the coverage and transferability of ByteFF.
@@ 3.2 Torsional Potential Energy Surfaces
To evaluate the accuracy of ByteFF in predicting torsional PES, and directly compare with the performance of other force fields, we performed comprehensive benchmarks using the TorsionNet500 and BDTorsion datasets. TorsionNet500 48 is a benchmark dataset consisting of 500 chemically diverse fragments relevant to biological and pharmaceutical applications. The BDTorsion dataset is curated in this work. It is divided into Non-Ring and In-Ring torsion subsets and each contains 1000 chemically diverse fragments. For both TorsionNet500 and BDTorsion datasets, dozens of conformations are sampled by scanning torsion angle, then QM-relaxed, and tagged with QM energy labels for each fragment. Here, two metrics were used to evaluate the accuracy of the predictions from various force fields with respect to the ground truth from QM: the RMSE and Boltzmann RMSE, with RMSE assessing the overall discrepancy while Boltzmann RMSE emphasizing the deviations near the energy minimum. When calculating the Boltzmann RMSE, we used 2.0 kcal/mol for both cutoff and scaling factors.
@@
As illustrated in Fig. 3, both ByteFF-gopt and ByteFF-joint significantly outperform competitors on all three benchmark datasets. On all the three datasets, ByteFF-gopt exhibits the best performance on both metrics, with the highest population of predictions with lower RMSE and Boltzmann RMSE. When tested on TorsionNet500, the Boltzmann RMSE of most predictions by ByteFF-gopt are within 0.5 kcal/mol, while none of the Boltzmann RMSE exceeds 1.17 kcal/mol, demonstrating the exceptional accuracy of ByteFF-gopt in terms of predicting torsional energy profiles. On the more challenging BDTorsion datasets, both RMSE and Boltzmann RMSE are generally higher for all the force fields being tested. Nevertheless, ByteFF-gopt and ByteFF-joint still outperformed competitors significantly, with most Boltzmann RMSE values below 1.0 kcal/mol (chemical accuracy). In these torsion profile benchmarks, the performance of ByteFF-joint is compromised by the addition of off-equilibrium energies and forces training data. This is inevitable due to the very limited fixed functional form of MMFFs.
@@
Figure 4: Example of the in-ring torsion prediction accuracy of various force fields. (a) 2D representation of the example molecule. (b) The comparison of torsional energy profiles predicted by various force fields and calculated from QM. (c) The zoomed-in comparison of torsional energy profiles predicted with torsion angles between -175 and -75.
@@
One key feature of organic molecules is the presence of a great variety of rings52, and in-ring torsion parameters have crucial impact on the conformational PES of rings. However, this is much less discussed in the development of MMFFs. In this work, we carefully constructed the in-ring torsion scan training dataset to improve the performance of ByteFF on in-ring torsions. We also curated the BDTorsion-InRing benchmark dataset, which consists of 1000 chemically diverse fragments with QM calculated in-ring torsion profiles. As shown in Fig. 3 (c) and (f), both ByteFF-gopt and ByteFF-joint significantly outperform competitors, with both RMSE and Boltzmann RMSE mostly better than chemical accuracy, respectively. As an example to illustrate the accuracy of ByteFF-gopt and ByteFF-joint in predicting in-ring torsional energy profile, we present a molecule from the BDTorsion-InRing dataset that contains a four-membered aliphatic ring (Fig. 4 (a)). On this molecule, we calculated the torsional energy profile for the in-ring torsion atoms highlighted in Fig. 4 (a), and compared PES predictions from various force fields with the QM references (Fig. 4 (b) and (c)). It is evident that both Espaloma-0.3.0 (green) and OpenFF-2.0.0 (blue) failed to predict the energy landscape in the vicinity of the local minimum, which could lead to incorrect predictions of ring conformations in geometric optimization or MD simulations. Though GAFF-2.2 captured the approximate shape of this torsional energy landscape, it overestimated the barrier near -130 while significantly underestimated the energy near -160, which led to significantly incorrect shifts of the predicted location of global energy minimum. In contrast, both ByteFF-gopt and ByteFF-joint predictions align well with QM references, which not only captured the shape of the energy landscape, but also accurately predicted the positions and energy differences of local minima.

### v2

@@ 2.1 Molecular Mechanics Force Field
$E^{\mathrm{MM}}=E_{\mathrm{bonded}}^{\mathrm{MM}}+E_{\mathrm{non-bonded}}^{\mathrm{MM}}$ (1)
$\begin{split}E_{\mathrm{bonded}}^{\mathrm{MM}}=&E_{\mathrm{bond}}^{\mathrm{MM}}+E_{\mathrm{angle}}^{\mathrm{MM}}+E_{\mathrm{proper}}^{\mathrm{MM}}+E_{\mathrm{improper}}^{\mathrm{MM}}\\ =&\sum_{\mathrm{bonds}}\frac{1}{2}k_{r,ij}(r_{ij}-r_{ij}^{0})^{2}+\sum_{\mathrm{angles}}\frac{1}{2}k_{\theta,ijk}(\theta_{ijk}-\theta_{ijk}^{0})^{2}\\ &+\sum_{\mathrm{propers}}\sum_{n_{\phi}}k_{\phi,ijkl}^{n_{\phi}}\left[1+\cos(n_{\phi}\phi_{ijkl}-\phi_{ijkl}^{n_{\phi},0})\right]\\ &+\sum_{\mathrm{impropers}}k_{\psi,ijkl}\left[1+\cos(2\psi_{ijkl}-\pi)\right]\\ E_{\mathrm{non-bonded}}^{\mathrm{MM}}=&E_{\mathrm{vdW}}^{\mathrm{MM}}+E_{\mathrm{Coulomb}}^{\mathrm{MM}}\\ =&\sum_{i<j}\left[4\epsilon_{ij}\left(\frac{\sigma_{ij}^{12}}{r_{ij}^{12}}-\frac{\sigma_{ij}^{6}}{r_{ij}^{6}}\right)+\frac{q_{i}q_{j}}{4\pi\epsilon_{0}r_{ij}}\right]\end{split}$ (2)
where the bond lengths $r$, angles $\theta$, proper torsion angles $\phi$ and improper torsion angles $\psi$ are rotationally invariant internal coordinates. The force field parameters include bonded parameters (equilibrium values $r^{0}$, $\theta^{0}$ and $\phi^{0}$, and force constants $k_{r}$, $k_{\theta}$, $k_{\phi}$ and $k_{\psi}$), and non-bonded parameters (van der Waals (vdW) parameters $\sigma$ and $\epsilon$, and partial charges $q$). A well-trained transferrable MMFF shall predict these parameters accurately for any given molecules. In this article, “torsion” refers to proper torsion unless otherwise specified. We also fix the phase angles $\phi_{ijkl}^{n_{\phi},0}$ at 0 for odd $n_{\phi}$ and $\pi$ for even $n_{\phi}$, and therefore ensure that the torsional energy is independent of the order being $ijkl$ or $lkji$.
@@
Additionally, there are two key philosophies worth considering when building general small molecule MMFFs: (1) Force field parameters should be dominated by local structures, so that the parameters trained from small molecules can be consistently transferred to similar structures in relatively large molecules. 9, 11 (2) Torsional energy profiles should be accurately captured, as the quality of torsion parameters significantly affects the conformational distribution of small molecules, thereby influencing the prediction of properties such as protein-ligand binding affinity. 32, 11, 33
@@ 2.2.2 Quantum chemistry methods and workflow
From the 2.4 million fragments, we created two QM datasets, namely the optimization dataset and the torsion dataset, at the B3LYP-D3(BJ)/DZVP level of theory. The same QM method is also used by OpenFF to generate their training data. 18, 19 This method achieves a good balance between accuracy (relative to CCSD(T)/CBS) and computational cost in recent benchmarks. 38, 39 For molecular conformational PES, more advanced methods such as $\omega$B97M-V are significantly more expensive but only marginally more accurate.
The optimization dataset was generated from the entire 2.4 million fragments. The 3D conformations of each fragment was initially generated by RDKit from its SMILES string, and then optimized using the geomeTRIC 40 optimizer at the chosen QM level. The relaxed structure is verified where no accidental bond breaking or formation happened during the structural relaxation. Finally, the Hessian matrix is calculated for each fragment using Q-Chem 6.141. The hessian matrix is further verified by checking all the eigenvalues (except for six near-zero values corresponding to translational and rotational modes) to ensure that a true local minimum is captured. More details about the workflow and the filtering criteria are provided in the supporting information.
The torsion dataset consists of two subsets, non-ring torsions and in-ring torsions, curated separately, comprising 2.2 million and 1.0 million frames respectively. For each non-ring torsion fragment, we first rotate the torsion angle in 15 increments from the optimized conformation, creating 24 initial frames that were then separately relaxed using the geomeTRIC optimizer with the rotated torsion angle constrained. While the in-ring torsions were scanned using a sequential frame-by-frame approach, with an early stopping strategy if the conformational energy is more than 20 kcal/mol higher than the unconstrained relaxed conformation, since those high-energy regimes are irrelevant in room-temperature molecular dynamics simulations. For both non-ring and in-ring subsets, the resultant conformers were also carefully filtered by the bond breaking or formation criteria. Due to the high computational cost of torsion scan, these calculations were performed with GPU4PySCF42, 43 for significant acceleration and cost reduction. The results were also verified to be consistent with Q-Chem.
@@ 2.3 ByteFF Model
Figure 1: Model structure of ByteFF. ByteFF predicts MMFF parameters in three steps. First, atom and bond features are extracted from the molecular graph and then projected into embeddings. Then, an edge-augmented graph transformer (EGT)44 is used to synergize the edge embeddings with the node-based attention mechanism. Lastly, the output module derives force field parameters while preserving the molecular symmetry and total charge.
ByteFF predicts both the bonded and non-bonded force field terms in one-pass. It leverages the state-of-the-art Edge-augmented Graph Transformer (EGT) 44 as its backbone to achieve accuracy and transferability. The output of EGT is further processed by a series of MLPs to derive force field parameters while preserving both the total charge and the physical symmetry of each force field term. The ML model structure is shown in Fig. 1 and consists of three major parts: featurization, GNN, and output modules. They are briefly described here, and more details are given in the supporting information.
Featurization module. For each molecule, we first convert the molecule structure into a graph, and detect chemically equivalent atoms and bonds with bond orders ignored. Atomic features are then detected from this graph, including the element type, formal charge, ring connectivity and minimum ring size. Additionally, to better describe the chemical environment of each atom, bond features including bond order and whether the bond is in ring, are also extracted. Replacing the discrete atom-typing rules used in conventional MMFFs, atom and bond features are projected into continuous vectorial embedding spaces, and concatenated into node and edge feature embeddings, respectively. Importantly, embeddings of chemically equivalent atoms and bonds are averaged before passed to the next module, which helps to preserve the chemical symmetry of the molecule.
GNN module. In the GNN module, a modified EGT architecture is employed to learn the chemical environment of molecules from both atom and bond features. This architecture synergizes the node and edge features in the attention mechanism to efficiently capture the structural information of the molecule. To ensure locality and speed up training and inference, the original global self-attention in EGT is modified such that attention is applied only to a local neighborhood of each node.
Output module. After a few layers of localized EGT, atom and bond hidden states generated by the GNN module are passed to the output module to predict each force field term individually. For each bonded term, the corresponding output module is designed to preserve the necessary symmetry, respectively. For the partial charge term, the output module is designed to preserve the total charge of the molecule following a bond-charge-correction (BCC) style postprocessing.
Additionally, an ensemble of five models randomly initialized is trained for improved predictive performance and uncertainty quantification.
@@ 2.4 Training Procedures
In the pre-training stage, the non-bonded parameters and force constants of proper torsions were fitted to GAFF-2.2 with mean squared error (MSE) losses, while the equilibrium values of bonded parameters were fitted with energy-based loss functions, using the optimization dataset. Meanwhile, the force constants of bonded terms in Eq. 2 were trained using a partial Hessian loss, evaluated by the mean absolute percentage error (MAPE) of Hessian blocks corresponding to bonds, angles and improper torsions, namely partial Hessian blocks. It has been reported that the partial Hessian blocks can be used to derive accurate force constants for single molecules. 45, 46, 47 Our design of the differentiable partial Hessian loss enables fitting force constants in batches, combining training accuracy with computational efficiency.
Given the significance of torsion profiles in the quality of force fields, in the training stage, we incorporated the curated torsion dataset to fit the force constants of proper torsions, while the other parameters were trained using optimization dataset following the same manner in the pre-training stage. Here, we employed the Boltzmann MSE loss functions 48 into an iterative optimization-and-training process to train the force constants of proper torsions using the torsion dataset. In each iteration, the QM-optimized geometries were refined by the force field, with the torsion angle constrained and atom positions restrained 18. The parameters were then trained on both the optimization and torsion datasets. As the force field’s accuracy improved, the positional restraint force constant was gradually reduced. Additionally, L1-norm regularization was applied to the force constants of proper torsions to restrain redundant degrees of freedom. The combination of pre-training and training stages yields the ByteFF-gopt model.
@@
Table 1: Summary of training settings for different stages
@@
Datasets and fitting targets optimization dataset: - GAFF-2.2 $\sigma/\epsilon/q/k_{\phi}$ $E^{\mathrm{MM}}_{\mathrm{bond}}$ / $E^{\mathrm{MM}}_{\mathrm{angle}}$ / $E^{\mathrm{MM}}_{\mathrm{improper}}$ QM Hessian matrix optimization dataset: - GAFF-2.2 $\sigma/\epsilon/q$ $E^{\mathrm{MM}}_{\mathrm{bond}}$ / $E^{\mathrm{MM}}_{\mathrm{angle}}$ / $E^{\mathrm{MM}}_{\mathrm{improper}}$ Hessian matrix torsion dataset: - QM energy - L1-norm of $k_{\phi}$ optimization dataset: same as training torsion dataset: same as training off-equilibrium dataset: - QM energy/force
@@ 2.5 Benchmark Datasets and Metrics
To be compatible with the Amber family of protein force fields, the non-bonded terms of ByteFF were trained to reproduce the AM1-BCC charges 49, 50 and GAFF-2.2 vdW parameters. Therefore, in this work we focus on the intramolecular PES performance relative to QM references of ByteFF, and leave the further improvement of non-bonded terms as a future work.
Three different types of benchmarks were used to provide a thorough assessment of ByteFF. First, the large-scale industrial collaborative OpenFFBenchmark public dataset 51 was used to quantify how well the force field-relaxed geometries and energies reproduced the QM-relaxed references, without constraints. Three different metrics were used, namely root-mean-square deviation of atomic positions (RMSD), torsion fingerprint deviation (TFD) 52, and relative energy difference ($\Delta\Delta E$) defined by OpenFF team.53 Second, the crucial torsional energy profile performance was benchmarked on two different datasets, namely the publicly available TorsionNet500 dataset 54 and the BDTorsion dataset curated in-house. For consistency, energies of conformations in TorsionNet500 54 dataset were re-calculated with the B3LYP-D3(BJ)/DZVP level of theory. On these datasets, the root mean squared error (RMSE) and Boltzmann-weighted RMSE of each torsional energy profile were calculated and compared.
Beyond the constrained (torsional energy) and unconstrained (OpenFFBenchmark) optimization benchmarks, the performance of ByteFF was further benchmarked on the QM references energies and forces calculated on the off-equilibrium dataset, which includes subsets of the SPICE dataset55 and the RNA Structure Atlas56. On this dataset, RMSE of energies and forces were calculated to quantify if the PES is properly captured beyond local minima.
Rigorous procedures were followed in the benchmarking of ByteFF models, which ensured no data leakage between the training and testing data. More details of these benchmarks can be found in supporting information.
@@ 3.1 Diversity of Various Datasets
Using the Morgan-based torsional fingerprint defined above, we quantify the diversity of SPICE, GEOM and BDTorsion datasets in terms of torsional local environment. Specifically, in all three datasets, only torsions with circular standard deviation 57 greater than 0.3 were included and used in fingerprint analysis. The diversity of each dataset was then visualized using the standard t-SNE (t-Distributed Stochastic Neighbor Embedding) algorithm. In Fig. 2, the scatter dots are colored by the element types of the two center atoms (the rotatable bond) in each torsion. The most frequent rotatable bond is C-C, followed by C-N rotatable bonds and other bonds including C-O, C-S, etc. It is evident that our torsion dataset excels in the diversity and comprehensiveness of its chemical space coverage, while the SPICE and GEOM datasets are more sparse. In comparison, both SPICE and GEOM datasets show obvious vacancies, corresponding to the absence of certain chemical torsion patterns. For example, the uncovered areas around (-45, -20) in Fig. 2 (a) and (b) correspond to in-ring torsions of protonated pyridine and imidazole, as well as their derivatives, which are indispensable for the training of force fields for bio-organic molecules. Such an expansive coverage of the chemical space in the training dataset provides a solid foundation for the coverage and transferability of ByteFF.
@@ 3.2 Torsional Potential Energy Surfaces
To evaluate the accuracy of ByteFF in predicting torsional PES, and directly compare with the performance of other force fields, we performed comprehensive benchmarks using the TorsionNet500 and BDTorsion datasets. TorsionNet500 54 is a benchmark dataset consisting of 500 chemically diverse fragments relevant to biological and pharmaceutical applications. The BDTorsion dataset is curated in this work. It is divided into Non-Ring and In-Ring torsion subsets and each contains 1000 chemically diverse fragments. For both TorsionNet500 and BDTorsion datasets, dozens of conformations are sampled by scanning torsion angle, then QM-relaxed, and tagged with QM energy labels for each fragment. Here, two metrics were used to evaluate the accuracy of the predictions from various force fields with respect to the ground truth from QM: the RMSE and Boltzmann RMSE, with RMSE assessing the overall discrepancy while Boltzmann RMSE emphasizing the deviations near the energy minimum. When calculating the Boltzmann RMSE, we used 2.0 kcal/mol for both cutoff and scaling factors.
@@
As illustrated in Fig. 3, both ByteFF-gopt and ByteFF-joint significantly outperform competitors on all three benchmark datasets. On all the three datasets, ByteFF-gopt exhibits the best performance on both metrics, with the highest population of predictions with lower RMSE and Boltzmann RMSE. When tested on TorsionNet500, the Boltzmann RMSE of most predictions by ByteFF-gopt are within 0.5 kcal/mol, while none of the Boltzmann RMSE exceeds 1.33 kcal/mol, demonstrating the exceptional accuracy of ByteFF-gopt in terms of predicting torsional energy profiles. On the more challenging BDTorsion datasets, both RMSE and Boltzmann RMSE are generally higher for all the force fields being tested. Nevertheless, ByteFF-gopt and ByteFF-joint still outperformed competitors significantly, with most Boltzmann RMSE values below 1.0 kcal/mol (chemical accuracy). In these torsion profile benchmarks, the performance of ByteFF-joint is compromised by the addition of off-equilibrium energies and forces training data. This is inevitable due to the very limited fixed functional form of MMFFs.
@@
Figure 4: Example of the in-ring and non-ring torsion prediction accuracy of various force fields. As examples to show the accuracy of ByteFF models in predicting torsional energy profiles, an in-ring (a-b) and a non-ring (c-d) example molecule are provided. The torsional energy profiles predicted by various force fields are compared with the QM references and shown for each example molecule.
@@
One key feature of organic molecules is the presence of a great variety of rings58, and in-ring torsion parameters have crucial impact on the conformational PES of rings. However, this is much less discussed in the development of MMFFs. In this work, we carefully constructed the in-ring torsion scan training dataset to improve the performance of ByteFF on in-ring torsions. As shown in Fig. 3 (c) and (f), both ByteFF-gopt and ByteFF-joint significantly outperform competitors, with both RMSE and Boltzmann RMSE mostly better than chemical accuracy, respectively. As an example to illustrate the accuracy of ByteFF-gopt and ByteFF-joint in predicting in-ring torsional energy profile, we present a molecule from the BDTorsion-InRing dataset that contains a four-membered aliphatic ring (Fig. 4 (a)). On this molecule, we calculated the torsional energy profile for the in-ring torsion atoms highlighted in Fig. 4 (a), and compared PES predictions from various force fields with the QM references (Fig. 4 (b)). It is evident that both Espaloma-0.3.0 (green) and OpenFF-2.0.0 (blue) failed to predict the energy landscape in the vicinity of the local minimum, which could lead to incorrect predictions of ring conformations in geometric optimization or MD simulations. Though GAFF-2.2 captured the approximate shape of this torsional energy landscape, it overestimated the barrier near -130 while significantly underestimated the energy near -160, which led to significantly incorrect shifts of the predicted location of global energy minimum. In contrast, both ByteFF-gopt and ByteFF-joint predictions align well with QM references, which not only captured the shape of the energy landscape, but also accurately predicted the positions and energy differences of local minima. In addition to accurately predicting in-ring torsional energy profiles, ByteFF models also excel in the prediction of non-ring torsions, as illustrated with the example molecule in Fig. 4 (c). As shown in Fig. 4 (d), ByteFF-gopt and ByteFF-joint successfully predicted the torsional energy landscape of the highlighted non-ring torsion, which is highly consistent with the QM reference, while other force fields failed to predict both the positions of energy minima and the height of the energy barrier. More example molecules are provided in Fig. S2-S3 including both in-ring and non-ring torsions, where ByteFF models predicted the torsional energy profiles with exceptional accuracy against the QM reference, while other force fields failed to achieve.